OpenStudy (anonymous):

Which function could be the parent of y = 3|x| ?

OpenStudy (anonymous):

a. y=|x| b. y=3x c. y=x^3 d. y=3+x I'm thinking that it's y=3x but I'm not very sure

OpenStudy (shubhamsrg):

I am confused about (a) and (b).

OpenStudy (shubhamsrg):

90% a and 10% b

OpenStudy (anonymous):

its a

OpenStudy (shubhamsrg):

Hmm, must be "a" only then since you say it! ;)

OpenStudy (anonymous):

ok i was stuck between the first two, thank you!

OpenStudy (anonymous):

Parent functions can be transformed to create other members in a family of graphs.

OpenStudy (shubhamsrg):

What inclined me to say it could have been b also since if you take modulus of 3x, it comes to be 3|x|

OpenStudy (anonymous):

but parent function ie originator graph is y=mod(x) as y=3x is a straight line......isnt it?

OpenStudy (anonymous):

@shubhamsrg

OpenStudy (shubhamsrg):

Yep, that reasoning sounds more sound. Hence b it should be.

OpenStudy (shubhamsrg):

I mean a, a it should be.